OpenStudy (anonymous):

Find the second derivative of y=((cos x)/(1+sin x))

OpenStudy (anonymous):

Use the quotient rule.

OpenStudy (anonymous):

Yea i knw . can i have a step by step

OpenStudy (anonymous):

Quotient rule is as follows:\[\left( \frac{ f(x) }{ g(x) } \right)'=\frac{ f'(x)g(x)-f(x)g'(x) }{ g^2(x) }| g(x) \ne 0\]

OpenStudy (anonymous):

So take the derivative of the top and multiply it by the bottom, then subtract the derivative of the bottom multiplied by the top. And all of this is over whatever is in the bottom squared.